# Autoregressive Stochastic Clock Jitter Compensation in Analog-to-Digital Converters

### v2

@@ III-B Kalman smoother
$\begin{cases}\xi_{n}=\varphi\xi_{n-1}+\epsilon_{n}\\ y_{n}=x_{n}+\xi_{n}x^{\prime}_{n}+w_{n},\par\end{cases}$ (18)
Our implementation of the Kalman filter and smoother routine is described in Algorithm 2. Even though it pretty much follows the standard textbook routine (see e.g. [31]), a slight modification was engineered to handle missing data at non pilots positions: the measurements covariances $R_{n}$ are there set to be =+absent=$\infty$= + ∞, making the Kalman gain null. This correctly encodes the fact that the actual tracked autoregressive process is a subsampled version of the original available through pilots; more concretely the Kalman filter will optimally predict the process $\xi_{n}=\varphi^{K+1}\xi_{n-K-1}+\nu_{n}$ with $\nu_{n}\sim\mathcal{N}(0,(1-\varphi^{2(K+1)})/(1-\varphi^{2}))$, and the smoother will interpolate between them using the assumed dynamics (18). Notice that the exact observation “matrices” $x_{n}^{\prime}$ are replaced with their estimations $y^{\prime}_{n}$. The flowchart 2 details out and clarifies our implementation.
@@ IV Simulation results
Our test signal $x$ is throughout this entire section a bandlimited Gaussian process (cfr. Section II) with cutoff frequency $40$ MHz, then sampled at $100$ MS/s and with normalized power so that $\sigma_{x}^{2}=1$. Tests with different bandwidths and sampling rates led us to conclude that their impact on the algorithms performances is negligible.
The digital signal is generated by applying the desired brickwall filter to a sequence $\mathbf{x}$ of $N=2^{18}$ independent real-valued Gaussian random variables; the samples of $\mathbf{x}$ are assumed to be taken at the time instants $nT_{s}$ and the jittered samples $x(nT_{s}+\xi_{n})$ can be obtained by processing the analog counterpart of $\mathbf{x}$ via the standard Whittaker-Shannon interpolation formula so that
@@
Moreover the training sequence of known pilot samples $x_{n}$ for $n\in I$ is time-multiplexed with the received signal samples. We chose a constant sequence of pilots that left the received signal’s variance unchanged. In a first batch of simulations we test the interplay between background white noise and jitter, and its effects on the proposed algorithms. In Figure 3 we keep constant (to two different levels) the amount of jitter. We sweep through $30$ different white noise power levels (and run $5$ experiments each) so that the NDR is between $-20$ and $10$ dB and the overall SINADR (cfr. formula (6)) is assessed. Two different pilot samples densities are considered too, while keeping fixed both the blocks length $C=500$ and the polynomial degree $d=4$ for Algorithm 1. The Kalman smoother has oracle knowledge of $\sigma_{w}^{2}$, $\sigma_{\epsilon}^{2}$ and $\varphi=0.999$. As expected, the Kalman smoother displays superior performances throughout the entire NDR range; improvements after compensation seem to be achievable up to $1-2$ dB, after which the jitter disturbance is essentially drowned in noise and thus no longer distinguishable (and leading).
@@
In Figure 5 we show the algorithms performances with increasing levels of jitter and constant noise power; the latter is fixed so that the NDR is $-10$ dB at the lowest jitter level. Compensated signals show $6-15$ dB improvements in SINADR up to $\approx 4\%$ jitter, where the Kalman smoother is taken over by the polynomial method that keeps delivering a consistent improvement up to $10\%$ of jitter. Interestingly, beyond a certain point the Kalman smoother’s performance degrades rapidly. This stems from how it assesses the trustworthiness of measurements via the ratio $\sigma_{\epsilon}^{2}/\sigma_{w}^{2}$, and from the fact that in high‑SNR, low‑NDR scenarios (as in Figure 5), the Kalman gain approaches unity - making the posterior essentially the raw measurement and thus highly sensitive to outliers. Of course white noise is one reason for outliers, but when the jitter is large and the NDR small (see Figure 6), there are other distortion sources that the Kalman scheme is not actively accounting for, most prominently the approximation $x^{\prime}_{n}\approx y^{\prime}_{n}$ and the increasingly weakening of the linear approximation model (4).
@@
Figure 6: Outliers-dominated landscape in a high-jitter-low-noise scenario.
@@ V Conclusions and future research
In the present paper we proposed two algorithms to address the stochastic distortion appearing in baseband signals produced by local oscillators in Analog-to-Digital Converters. A mathematically clean and convenient description of the jitter noise as autoregressive process of order $1$ was also reiterated. The Kalman smoother approach shows robust and nearly optimal performances in a wide range of scenarios (cfr. Figure 3), and both methods display significant improvements in SINADR levels even with very low sample pilots densities (Figure 4). Their impact is assessed also when pushing the validity of the linear approximation (4) to a breaking point, and we discovered that for very large jitter the polynomial compensation algorithm could be more suitable, as less sensitive to outliers (Figure 5). Future research will focus on assessing the robustness of the Kalman smoother in the face of parameter inaccuracies, on incorporating additional distortions (e.g. phase noise when the signal is upconverted to RF or strong nearby blockers already contaminated with Digital-to-Analog Converter (DAC) jitter) in the compensation routines as well as on designing optimal training sequences of pilot samples.

### v3

@@ III-B Kalman smoother
$\begin{cases}\xi_{n}=\varphi\xi_{n-1}+\epsilon_{n}\\ y_{n}=x_{n}+\xi_{n}x^{\prime}_{n}+w_{n},\end{cases}$ (18)
@@
Our implementation of the Kalman filter and smoother routine is described in Algorithm 2. Even though it pretty much follows the standard textbook routine (see e.g. [37]), a slight modification was engineered to handle missing data at non pilots positions: the measurements covariances $R_{n}$ are there set to be =+absent=$\infty$= + ∞, making the Kalman gain null. This correctly encodes the fact that the actual tracked autoregressive process is a subsampled version of the original available through pilots; more concretely the Kalman filter will optimally predict the process $\xi_{n}=\varphi^{K+1}\xi_{n-K-1}+\nu_{n}$ with $\nu_{n}\sim\mathcal{N}(0,(1-\varphi^{2(K+1)})\sigma_{\epsilon}^{2}/(1-\varphi^{2}))$, and the smoother will interpolate between them using the assumed dynamics (18). Notice that the exact observation “matrices” $x_{n}^{\prime}$ are replaced with their estimations $y^{\prime}_{n}$. The flowchart 2 details out and clarifies our implementation.
III-C Complexity analysis
The Kalman smoother processes each sample within the stream; a single iteration in both forward and backward passes has constant cost, thus the total cost for $N$ samples is $O(N)$.
In the polynomial technique, the matrix to be inverted in (15) is $d\times d$; the full cost of (15), including matrix multiplications, is thus $O(Cd^{2}+d^{3})$. Since this is repeated $L$ times, the total cost for calculating the polynomial coefficients will be $O(L(Cd^{2}+d^{3}))$; the polynomial evaluations cost $O(Nd)$, bringing the total complexity to $O(Nd+L(Cd^{2}+d^{3}))$.
@@ IV Simulation results
Our test signal $x$ is throughout most of this section is a bandlimited Gaussian process (cfr. Section II) with cutoff frequency $40$ MHz, then sampled at $100$ MS/s and with normalized power so that $\sigma_{x}^{2}=1$. The digital signal is generated by applying the desired brickwall filter to a sequence $\mathbf{x}$ of $N=2^{18}$ independent real-valued Gaussian random variables; the samples of $\mathbf{x}$ are assumed to be taken at the time instants $nT_{s}$ and the jittered samples $x(nT_{s}+\xi_{n})$ can be obtained by processing the analog counterpart of $\mathbf{x}$ via the standard Whittaker-Shannon interpolation formula so that
@@
Moreover the sequence of known pilot samples $x_{n}$ for $n\in I$ is time-multiplexed with the received signal samples. For the experiments within this section we chose a constant sequence of pilot samples that left unchanged the signal’s variance. In a first batch of simulations we test the interplay between background white noise and jitter, and its effects on the proposed algorithms. In Figure 4 we keep constant (to two different levels) the amount of jitter. We sweep through $30$ different white noise power levels (and run $5$ experiments each) so that the NDR is between $-20$ and $10$ dB and the overall SINADR (cfr. formula (6)) is assessed. Two different pilot samples densities are considered too, while keeping fixed both the blocks length $C=500$ and the polynomial degree $d=4$ for Algorithm 1. The Kalman smoother has oracle knowledge of $\sigma_{w}^{2}$, $\sigma_{\epsilon}^{2}$ and $\varphi=0.999$. As expected, the Kalman smoother displays superior performances throughout the entire NDR range; improvements after compensation seem to be achievable up to $1-2$ dB, after which the jitter disturbance is essentially drowned in noise and thus no longer distinguishable (and leading).
@@
In Figure 6 we show the algorithms performances with increasing levels of jitter and constant noise power; the latter is fixed so that the NDR is $-10$ dB at the lowest jitter level. Compensated signals show $6-15$ dB improvements in SINADR up to $\approx 4\%$ jitter, where the Kalman smoother is taken over by the polynomial method that keeps delivering a consistent improvement up to $10\%$ of jitter. Interestingly, beyond a certain point the Kalman smoother’s performance degrades rapidly. This stems from how it assesses the trustworthiness of measurements via the ratio $\sigma_{\epsilon}^{2}/\sigma_{w}^{2}$, and from the fact that in high‑SNR, low‑NDR scenarios (as in Figure 6), the Kalman gain approaches unity - making the posterior essentially the raw measurement and thus highly sensitive to outliers. Of course white noise is one reason for outliers, but when the jitter is large and the NDR small, there are other distortion sources that the Kalman scheme is not actively accounting for, most prominently the approximation $x^{\prime}_{n}\approx y^{\prime}_{n}$ and the increasingly weakening of the linear approximation model (4).
We note that the parameters $d$, $C$ and the pilot density interact in the polynomial technique. In these experiments, we did not attempt to identify their optimal combination; instead, we chose sub-optimal values empirically to avoid both over-fitting and under-fitting. The problem of optimally interpolate Wiener-type processes with polynomial functions has been already studied in e.g. [6]. Thus a more rigorous investigation of this interplay is left to future work.
A last numerical simulation is shown in Figure 7, where we sweep through different signal bandwidths to show the SINADR of the compensated signals, in a similar fashion as in Figure 4; as the bandwidth increases we kept either the NDR (left) or the SNR (right) constant.
Figure 7: Algorithm performances vs signal bandwidth. Fixed NDR $=-5$ dB (left) and fixed SNR $=40$ dB (right). $\varphi=0.999$, $\sigma_{\xi}/T_{s}=1.5\cdot 10^{-2}$.
IV-A Suppression of jitter from blockers
The numerical examples displayed in the previous section deal with jitter suppression seen as in-band distortion; the power spectral densities of compensated and uncompensated signals look mostly the same because the term $\xi x^{\prime}$ is much weaker (in variance) than the signal itself if (II-B) holds; this can be rigorously proved by modifying the proof of Proposition II.2. The fact that the term $\xi x^{\prime}$ produces mostly in-band distortions is not entirely obvious, because its PSD is not compactly supported. However its out-of-band contributions are negligible; this can be seen by studying the integral function
[…]
that has two peaks at $|\nu|\approx\overline{\omega}$ and sharply declines to 00 as soon as $|\nu|>\overline{\omega}$.
To be better able to visually show the spectral impact of the proposed methods, we outline a different (and somewhat, in this section, ideal) scenario borrowed from radar applications; we assume a bistatic pair consisting of an illuminator of opportunity (IO) at known location and an antenna receiver with two channels, one for surveillance (SC) and one for reference (RC). The RC is responsible for acquiring high fidelity “copies” of the IO waveform, while the SC is primarily responsible for capturing targets’ echo. The SC is however also corrupted by Direct-Path Interferences (DPI) which are usually much stronger than echoes, and can thus entirely overshadow them.
The signal after ADC sampling can be written as
$z_{n}=x_{n}+\xi_{n}x_{n}^{\prime}+\zeta_{n}+\xi_{n}\zeta_{n}^{\prime}+w_{n}$ (21)
where $\zeta_{n}$ is the contribution from the DPI while $x_{n}$ is the wanted signal, and $\zeta_{n}+x_{n}$ are assumed to be known if $n\in I$; by hypothesis $\sigma_{\zeta}^{2}\gg\sigma^{2}_{x}$ and $\sigma_{\zeta^{\prime}\xi}^{2}>\sigma^{2}_{x}$ so that even the jitter coming from the blocker overpowers the signal of interest.
Figure 8: PSD at different compensation stages.
We picture a situation in which the term $\zeta_{n}$ can be entirely canceled with negligible impact on the others (at least if the target is moving and there is a nontrivial doppler component in $x_{n}$), for example projecting it away using subspace methods in combination with the RC signal knowledge, see e.g. [10]. After de-jittering $z_{n}$ using the methods presented in the previous section and optimally interpolating the “holes” left after removing the pilot samples using the Gerchberg-Papoulis algorithm, we then project away $\zeta_{n}$ and what is left is the sought echo signal. In Figure 8 are shown examples of power spectral densities of the different compensation stages described in this section.
@@ V Conclusions and future research
In the present paper we proposed two algorithms to address the stochastic distortion appearing in baseband signals produced by local oscillators in Analog-to-Digital Converters. A mathematically clean and convenient description of the jitter noise as autoregressive process of order $1$ was also reiterated. The Kalman smoother approach shows robust and nearly optimal performances in a wide range of scenarios (cfr. Figure 4), and both methods display significant improvements in SINADR levels even with very low sample pilots densities (Figure 5). Their impact is assessed also when pushing the validity of the linear approximation (4) to a breaking point, and we discovered that for very large jitter the polynomial compensation algorithm could be more suitable, as less sensitive to outliers (Figure 6). Future research will focus on assessing the robustness of the Kalman smoother in the face of parameter inaccuracies, on incorporating additional distortions (e.g. phase noise when the signal is upconverted to radio frequency, strong nearby blockers already contaminated with Digital-to-Analog Converter (DAC) jitter, power amplifiers nonlinearities, harmonic distortions etc.) in the compensation routines as well as on designing optimal training sequences of pilot samples.